OpenStudy (anonymous):

2y+8x=24 (1,-2) and (0,4)

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

that would help

OpenStudy (anonymous):

How do I break this down?

OpenStudy (anonymous):

2y+8x=24 (1,-2) and (0,4)

OpenStudy (anonymous):

are you trying to prove/disprove the point, i am not sure what you mean by "break it down"

OpenStudy (anonymous):

... is there a solution the the equation?

OpenStudy (anonymous):

solving for x or y will give 0 as an answer

OpenStudy (anonymous):

2y+8x=24 2y=24-8x 1/2[2y=24-8x] y=12-4x sub this into original equation 2(12-4x)+8x=24 24-8x-8x=24 0=0

OpenStudy (anonymous):

same if you also solve for x and substitute it back in

OpenStudy (anonymous):

2y+8x=24 1/2[2y+8x=24] y+4x=12 1/4[4x=12-y] x=12-y/4 sub this in to original equation 2y+8(12-y/4)=24 2y+2(12-y)=24 2y+24-2y=24 0=0